OpenStudy (anonymous):

Evaluate the limit of x/(absolute value of x) as x approaches 0- (from the left side). ???

zepdrix (zepdrix):

\[\Large \left(|x|\right)' \quad=\quad \frac{x}{|x|}\]

OpenStudy (anonymous):

\[ f(x) = \frac x {|x|} \]

OpenStudy (anonymous):

Since \(x\leq 0\) we know \(|x| = -x\)\[ \frac x{|x|}\to \frac x{-x} \to -1 \]

OpenStudy (anonymous):

But if x\[\le0\], shouldn't that -x be \[\left| -x \right|\]?

OpenStudy (anonymous):

Suppose \(x=-1<0\) Then \(|x|=|-1| = 1 = -x\)

OpenStudy (anonymous):

I see what you're saying now.

OpenStudy (anonymous):

The thing is that \(-x\) doesn't make \(x\) negative. It changes the sign of \(x\). Negative \(x\) values need to change sign to become positive.